Today hector is 22 years older than his sister. In 5 years he will be 3 times as old as she will be. How old is hector and his sister today?
Answer by macston(5194)

H=hector's age; S=sister's age
H=S+22yrs
H+5=3(S+5) Substitute for H.
S+22+5=3S+15
S+27=3S+15 Subtract 15 from each side.
S+12=3S Subtract S from each side.
12=2S Divide each side by 2.
6=S ANSWER 1: Sister is 6 years old
H=S+22yrs=6yrs+22yrs=28 yrs ANSWER 2: Hector is 28 years old.
CHECK:
H+5=3(S+5)
28yrs+5yrs=3(6yrs+5yrs)
33yrs=3(11yrs)
33yrs=33yrs